solve

In a potassium chloride unit, the distance between the potassiumion (K+) and the chloride ion (Cl-) is2.80 10-10 m.(a) Calculate the energy required to separatethe two ions to an infinite distance apart. (Model the two ions astwo point particles initially at rest.)
1 eV

(b) If twice the energy determined in Part (a) is actuallysupplied, what is the total amount of kinetic energy that the twoions have when they were an infinitedistance apart?
2 eV
Answer #1
in the given problem we get thatWext = ΔK + ΔU= 0 - Ui=- (k q- q+) / r=- k (- e) e / r= k e2 / r= (9 x 109 N . m2 / C2) (1.602 x10-19C)2 / (2.80 x10-10 m)= ......... J= (.......... J) (1 eV / 1.602 x 10-19 J)= ......... eV
answered by: anonymous person